I have analyzed the data from the dark fringe offset reduction to measure the contrast defect.
Figure 1 shows the signal of interest (B1 power, ITF optical gain, and RF DARM error signal)
Figure 2 shows looking at the relation between optical gain and B1 power and offset of 130uW seem to be present on B1 PD1.
Figure 3 shows the comparion between the RF DARM error signal and the B1 power, the relation here is also quadratic as the RF DARM error signal should be proportional to the DARM offset in meter (and the B1 power is quadratic with the DARM offset in meters). This yields an offset of 120uW.
After the unlock I could check that the PD itself has an offset of -7uW, so the numbers above correspond to actually 137uW and 127uW. Given that B1 PD1 sees only half of the power, the total contrast defect light is about 260uW. This is much higher than the ~120uW we had before increase of input power during the October commissioning break